# Multi-view X-ray Image Synthesis with Multiple Domain Disentanglement from CT Scans

### v1

@@ Abstract.
Generative Adversarial Networks, Image Synthesis, X-ray, Style Disentanglement, Multi-domains
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; XX XX - XX XX, XXXX; XX, XX.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Appearance and texture representations
@@ 2.3. GAN-based Image synthesis
In recent years, generative adversarial networks (GANs) have achie-ved remarkable success in image synthesis (Zhang et al., 2018b), image translation (Isola et al., 2017), and image editing (Alaluf et al., 2021; Zhou et al., 2023). Building upon progressive GAN (Karras et al., 2017), Karras et al. proposed StyleGAN (Karras et al., 2019), which enhanced the quality of the generated image and allowed the network to decouple different features, providing a more controllable synthesis strategy. StyleGAN2 (Karras et al., 2020) redesigned the instance normalization scheme to remove the water droplet-like artefacts existing in StyleGAN, leading to higher-quality outputs. Some methods have attempted to control generated images by exploring the latent space of GANs (Collins et al., 2020; Shen et al., 2020; Roich et al., 2022; Zhou et al., 2022). Despite their remarkable process, most deviations of GANs still concentrate on data augmentation in medical imaging, with limited research conducted on synthesizing cross-dimension medical images (Iglesias et al., 2023). One of the primary challenges inhibiting GANs from the CT2X task is the synthesis of multi-view results, as GANs have limited awareness of pose information (Iglesias et al., 2023; Liao et al., 2024).
@@ 3.3. Style Decoupling Encoder
Figure 3. An illustration of consistency regularization. $f_{s}^{X}$ and $f_{s}^{DRR}$ represent the style code extracted from X-ray and DRR images. $f_{s}^{\hat{X}}$ and $f_{s}^{\hat{DRR}}$ are the style features of synthesised results using X-ray and DRR, respectively. $f_{c}^{\hat{X}}$ and $f_{c}^{\hat{DRR}}$ share the same contents.
@@ Acknowledgements.
This work was supported in part by grants from National Key Research and Development Program of China (2023YFB4706003), National Natural Science Foundation of China (62176268), and Beijing Natural Science FoundationJoint Funds of Haidian Original Innovation Project (L232022).

### v2

@@ Abstract.
Generative Adversarial Networks; Image Synthesis; X-ray; Style Disentanglement; Multi-domains
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681154isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Appearance and texture representations
@@ 1.4. GAN-based Image synthesis
In recent years, generative adversarial networks (GANs) have achie-ved remarkable success in image synthesis (Zhang et al., 2018b), image translation (Isola et al., 2017), and image editing (Alaluf et al., 2021). Building upon progressive GAN (Karras et al., 2017), Karras et al. proposed StyleGAN (Karras et al., 2019), which enhanced the quality of the generated image and allowed the network to decouple different features, providing a more controllable synthesis strategy. StyleGAN2 (Karras et al., 2020) redesigned the instance normalization scheme to remove the water droplet-like artefacts existing in StyleGAN, leading to higher-quality outputs. Some methods have attempted to control generated images by exploring the latent space of GANs (Collins et al., 2020; Shen et al., 2020; Roich et al., 2022). Despite their remarkable process, most deviations of GANs still concentrate on data augmentation in medical imaging, with limited research conducted on synthesizing cross-dimension medical images (Iglesias et al., 2023). One of the primary challenges inhibiting GANs from the CT2X task is the synthesis of multi-view results, as GANs have limited awareness of pose information (Iglesias et al., 2023; Liao et al., 2024).
@@ 2.3. Style Decoupling Encoder
Figure 3. An illustration of consistency regularization. $f_{sty}^{X}$ and $f_{sty}^{DRR}$ represent the style code extracted from X-ray and DRR images. $f_{sty}^{\hat{X}}$ and $f_{sty}^{\hat{DRR}}$ are the style features of synthesised results using X-ray and DRR, respectively. $f_{c}^{\hat{X}}$ and $f_{c}^{\hat{DRR}}$ share the same contents.
@@ Acknowledgements.
This work was supported in part by grants from National Key Research and Development Program of China (2023YFC2508800), National Natural Science Foundation of China (62176268), Beijing Natural Science Foundation-Joint Funds of Haidian Original Innovation Project (L232022) and Fundamental Research Funds for the Central Universities (FRF-TP-24-022A). Thanks to OpenBayes.com for providing model training and computing resources.